# Surface solar radiation: AI satellite retrieval can outperform Heliosat and generalizes well to other climate zones

### v1

@@ Abstract
Accurate estimates of surface solar irradiance (SSI) are essential for solar resource assessments and solar energy forecasts in grid integration and building control applications. SSI estimates for spatially extended regions can be retrieved from geostationary satellites such as Meteosat. Traditional SSI satellite retrievals like Heliosat rely on physical radiative transfer modelling. We introduce the first machine-learning-based satellite retrieval for instantaneous SSI and demonstrate its capability to provide accurate and generalizable SSI estimates across Europe. Our deep learning retrieval provides near real-time SSI estimates based on data-driven emulation of Heliosat and fine-tuning on pyranometer networks. By including SSI from ground stations, our SSI retrieval model can outperform Heliosat accuracy and generalize well to regions with other climates and surface albedos in cloudy conditions (clear-sky index ¡ 0.8). We also show that the SSI retrieved from Heliosat exhibits large biases in mountain regions, and that training and fine-tuning our retrieval models on SSI data from ground stations strongly reduces these biases, outperforming Heliosat. Furthermore, we quantify the relative importance of the Meteosat channels and other predictor variables like solar zenith angle for the accuracy of our deep learning SSI retrieval model in different cloud conditions. We find that in cloudy conditions multiple near-infrared and infrared channels enhance the performance. Our results can facilitate the development of more accurate satellite retrieval models of surface solar irradiance.
@@
Solar radiation , surface solar irradiance , satellite retrieval , deep learning , emulation , Heliosat , Meteosat[tudelft]organization=Geosciences and Remote Sensing, Delft University of Technology,addressline=Stevinweg 1, city=Delft, postcode=2628 CN, state=South-Holland, country=The Netherlands
[bern]organization= IDAS, Bern University of Applied Sciences,addressline=Quellgasse 21, city=Biel, postcode=2502, state=Bern, country=Switzerland
Deep learning satellite retrieval of instantaneous surface solar radiation
Improves Heliosat retrieval accuracy by fine-tuning on ground station data
Can outperform Heliosat in cloudy conditions (clear-sky index ¡ 0.8)
Generalizes to remote locations and surface characteristics it was not trained on
We quantify the relevance of Meteosat channels for solar radiation retrieval
1 Introduction
Surface radiation estimates are required in solar resource assessments [1, 2, 3], solar forecast models [4, 5] and climate studies. High temporal and spatial coverage and resolution are essential for forecasting solar radiation for minutes to hours ahead (solar nowcasting) across large regions. While plant-level forecasts often rely on ground-based sensors [6, 7, 8], forecasts across spatially extended regions require surface radiation retrieved from geostationary satellites [9, 5, 10, 11, 12, 13]. Grid operators and energy providers use solar nowcasts to anticipate volatile solar energy and ensure grid stability. The relation between surface solar irradiance (SSI) and top-of-atmosphere reflectances measured from geostationary satellites has been explored and leveraged for retrieving SSI estimates for more than five decades [14, 15]. Satellite retrievals of SSI utilize top-of-atmosphere visible and infrared reflectances to estimate the amount of solar radiation that reaches the Earth’s surface.
Data-driven solar retrieval models can outperform physical models for single ground stations [16]. Deep learning has enabled increases in SSI retrieval accuracy of up to 20% compared to state-of-the-art retrievals [17, 18, 19, 20]. Recent data-driven retrieval models have been trained on pyranometer ground stations with geostationarily measured radiances as predictors to retrieve SSI by either post-processing previous SSI retrievals [21, 22] or by directly estimating SSI retrievals for individual ground stations [20, 23, 16, 17, 22]. SSI retrievals are expected to not only achieve high accuracy and low biases within the geographic region of their training set ground stations but also well beyond. In particular, an SSI retrieval should also be accurate in regions with other surface albedos and other climates, which have not been included in the development of the SSI retrieval model (out-of-domain generalization) [22, 20, 23, 24]. Existing machine-learning based SSI retrievals were found to lack generalizability and to feature artefacts showing up in SSI estimates at locations outside their training set domains [20, 24].
@@ 1 Introduction
The structure of this paper is outlined as follows. Section 2 reviews existing SSI retrieval methods. Section 3 introduces the datasets utilized in this study. The deep learning SSI retrieval and its development are detailed in section 4. Results of this study are discussed in section 5 and conclusions are provided in section 6.
@@ 2 Previous work
One of the first SSI satellite retrieval algorithms, Heliosat-1, included an empirical formula to calculate SSI [25]. An empirical clear-sky model estimated SSI in the absence of clouds and assuming a given atmospheric composition. The transmissivity of the clouds was then estimated by the ratio between top-of-atmosphere reflectivity measured and the minimum reflectivity over a month for cloud-free skies. The clear-sky reflectance mainly depends on the local solar zenith angle (SZA) and the surface albedo at the monitored site. Reliable surface albedo estimates are considered essential for accurate SSI retrievals [26]. Current state-of-the-art surface solar irradiance retrieval algorithms, such as CAMS [27], SARAH-3 [28], MSG-CPP [29, 30], and SolarGIS [31], use a combination of radiative transfer model (RTM) and the clear-sky index (CSI) to estimate SSI. The CSI is defined as the ratio between all-sky SSI and clear-sky SSI, , where the clear-sky SSI is calculated using a clear-sky RTM. The latter performs corrections to the clear-sky SSI due to attenuation from aerosols and trace gases. CAMS, SARAH-3 and MSG-CPP apply a cloudy-sky RTM to extract all-sky SSI based on cloud optical properties. RTM simulations are computationally expensive, so all three methods utilize look-up tables created by calculating radiative properties for various cloud and aerosol conditions. The observed cloud optical properties, aerosol properties and surface albedo are interpolated within the conditions provided by the look-up table to estimate global surface solar radiation. In MSG-CPP, the cloud optical properties are derived from near-infrared, visible light and surface albedo as parameters to a cloudy RTM look-up table [29, 32]. While the SSI retrieval algorithms differ, they are all based on the visible and partially near-infrared reflectivity from Meteosat Second Generation satellites.
@@
Recent research on deep learning (DL) for surface solar irradiance retrieval can be categorized into two main branches: post-processing of SSI derived from radiative transfer model and direct DL retrieval models. In post-processing, DL is employed as a post-processing step for existing retrieval algorithms by using spectral imagery and derived variables as input to improve on the initial SSI retrieval estimate. Cornejo-Bueno et al. [16] benchmarked a multitude of DL networks for post-processing Heliosat-2 and CAMS output, and found moderate improvements over a single station in Toledo, Spain. Verbois et al. [21] trained an extreme gradient boosting network to improve SSI estimates from HelioClim-3 [33] based on SSI from a group of 283 pyranometers spread across France. These methods have not been validated with regard to their generalizability to other regions outside the training set domain. Direct retrieval DL models, on the other hand, estimate SSI directly from visible or infrared satellite channels or ground station data. Quesada-Ruiz et al. [17] trained three model ensembles of ten fully connected feed-forward neural networks each on all visible and infrared channels pixel values of the Spinning Enhanced Visible and Infrared Imager (SEVIRI) plus a clear-sky SSI estimate and used SSI from eight ground stations in Europe as predictands (ground truth). Each ensemble was designed to estimate hourly SSI for different weather conditions: clear-sky (CSI 0.8), intermediate (0.8 ¿ CSI 0.4), and overcast conditions (CSI ¡ 0.4). A separate ensemble network of the same number of models predicted the clear-sky index itself. The method achieved an average RMSE skill score improvement of 17% compared to the Heliosat-2 method [34] across 28 validation stations. In another study, Jiang et al. [23] applied a convolutional residual network to estimate hourly SSI based on the Multifunctional Transport Satellites (MTSAT) series with 16x16 pixel patches of visible reflectances. The network was trained on 90 pyranometer stations spread out over different land surface classes and climates in China and inferred over the entirety of China. The network performance on eight validation pyranometers within the region showed an average RMSE of 93 . They did not compare those results with an existing radiative transfer based retrieval algorithm. The generalizability of retrieval method to out-of-sample geographical locations was not investigated by the authors.
@@
Verbois et al. [20] were the first to explore the out-of-sample generalizability of DL models for SSI retrievals. They trained a fully connected neural network on the hourly SSI from pyranometers over France incorporating 13 time steps of 15 minutes patches of 3x3 from three SEVIRI channels: the visible and (VIS006, VIS008) and infrared (IR108). They showed the performance of the DL network to be better than CAMS with an average RMSE skill score of 20% within the training domain of southeastern France, but outside this limited domain, in the validation set over the rest of France, performance dropped off significantly to skill scores of -100% to -40% for some stations. The worse performing stations were located in coastal regions with a generally lower surface albedo which were not represented in the training dataset. Verbois et al. [20] analyzed if average surface albedo impacted the generalization of the DL network and found a significant Spearman correlation of 0.346 (p¡0.01) between surface albedo and RMSE skill score. The results of Verbois et al. [20] show challenges in generalising a DL retrieval to unseen locations far differentiated from the training set.
@@ 3.1 Meteosat
Visible and infrared radiances from Earth that are measured by geostationary satellites provide the basis of SSI retrieval products. The Meteosat Second Generation (MSG) satellite at E, N is monitoring the regions of interest in our study – Europe and North Africa – with the Spinning Enhanced Visible and Infrared Imager (SEVIRI) that scans from East to West by rotating its sensor’s view. After each rotation, the sensor resets and scans the next line from North to South. One line scan is close to an instantaneous capture of longitudinal pixels, but individual scan lines have a delay associated with them, so an entire scan of the satellite’s disk takes approximately 12 minutes [35]. The ground resolution of the sensor is 3x3 km at nadir. All eleven visible and infrared channels are included as input to the emulator model training. The purpose of the channels can be summarized as follows [35]: The visible bands VIS006 and VIS008 are crucial for cloud detection, tracking, aerosol and land surface/vegetation monitoring. The near-infrared channel (IR016) helps distinguish snow from clouds and ice from water clouds, and provides information on aerosols. Infrared (IR039) primarily helps detect low clouds and fog, measure land/sea surface temperatures at night, and enhance low-level wind coverage through cloud tracking. The water vapor bands and (WV062, WV073) are meant for water vapor observation, using two channels at different tropospheric levels, aiding in thin cloud height allocation. Band (IR087) provides data on thin cirrus clouds and helps differentiate between ice and water clouds. Band (IR097) is utilized for tracking ozone patterns and contributing to numerical weather prediction by monitoring the total ozone field. Infrared and (IR108, IR120) are split window channels essential for measuring sea/land surface and cloud-top temperatures, and detecting cirrus and volcanic ash clouds. The last band (IR134) is a CO2 absorption channel and improves the height allocation of cirrus clouds.
3.2 Heliosat
Heliosat SARAH-3 Pfeifroth et al. [28] provides Meteosat-derived estimates of surface solar irradiance and other atmospheric variables related to clouds and radiation. It spans Meteosat First Generation (1977-2017) and MSG (2002-present) satellite retrievals from their respective instruments MVIRI and SEVIRI. The MVIRI is the predecessor of SEVIRI. SARAH-3 employs only the visible channels to keep the SARAH-3 record homogeneous in time.
The SARAH-3 algorithm draws on the cloud index method of the original Heliosat-1 algorithm by Cano et al. [25] to derive an estimated effective cloud albedo CAL. The cloud index method calculates the ratio between measured broadband reflectance , clear-sky reflectance and maximum reflectance as CAL . The clear-sky and maximum reflectance are estimated from the frequency distribution within a month, where the maximum reflectance is taken as the quantile in the month and the clear-sky reflectance as a stationary variate of lowest values. The lowest reflectance in the month is not taken as clear-sky reflectance due to cloud shadows which give low reflectance outliers in the tail of the distribution. The effective cloud albedo is converted to a cloud index with a smooth piecewise function defined in Dagestad [36]. The observed pixel-wise cloud index serves as input to a radiative transfer look-up table to retrieve SSI [32]. Instantaneous SSI from SARAH-3 is published with a two-week delay at a frequency of 30 minutes, so every second SEVIRI capture.

### v2

# Surface solar radiation: AI satellite retrieval can outperform Heliosat and generalizes to other climate zones

@@ Abstract
Accurate estimates of surface solar irradiance (SSI) are essential for solar resource assessments and solar energy forecasts in grid integration and building control applications. SSI estimates for spatially extended regions can be retrieved from geostationary satellites such as Meteosat. Traditional SSI satellite retrievals like Heliosat rely on physical radiative transfer modelling. We introduce a machine-learning-based satellite retrieval for instantaneous SSI and demonstrate its capability to provide accurate and generalizable SSI estimates across Europe. Our deep learning retrieval provides near real-time SSI estimates based on data-driven emulation of Heliosat and fine-tuning on pyranometer networks. By including SSI from ground stations, our SSI retrieval model can outperform Heliosat accuracy and generalize well to regions with other climates and surface albedos in cloudy conditions (clear-sky index 0.8). Our results indicate that the generalizability of a data-driven SSI retrieval model is not only related to the model training data or training method, but also depends on the amount of cloudiness present in the location at which SSI is retrieved with the data-driven model. We found that, in cloudy conditions, a model trained only on ground stations can estimate SSI accurately even in locations with different surface albedos, far away from the training test domain. We also show that the SSI retrieved from Heliosat exhibits large biases in mountain regions, and that training and fine-tuning our retrieval models on SSI data from ground stations strongly reduces these biases, outperforming Heliosat. Furthermore, we quantify the relative importance of the Meteosat channels and other predictor variables like solar zenith angle for the accuracy of our deep learning SSI retrieval model in different cloud conditions. We find that in cloudy conditions multiple near-infrared and infrared channels enhance the performance. Our results can facilitate the development of more accurate satellite retrieval models of surface solar irradiance.
@@
Solar radiation; surface solar irradiance; satellite retrieval; deep learning; emulation; Heliosat; Meteosat1 Introduction
Surface radiation estimates are required in solar resource assessments (Yang et al. 2022b; Carpentieri et al. 2023; Vuilleumier et al. 2020), solar forecast models (Yang et al. 2022a; Carpentieri et al. 2023) and climate studies. High temporal and spatial coverage and resolution are essential for forecasting solar radiation for minutes to hours ahead (solar nowcasting) across large regions. While plant-level forecasts often rely on ground-based sensors (Nouri et al. 2019; Kuhn et al. 2018; Barry et al. 2023), forecasts across spatially extended regions require surface radiation retrieved from geostationary satellites (Kosmopoulos et al. 2024; Carpentieri et al. 2025; Knol et al. 2021; Ayet and Tandeo 2018; Kellerhals et al. 2022). Grid operators and energy providers use solar nowcasts to anticipate volatile solar energy and ensure grid stability. The relation between surface solar irradiance (SSI) and top-of-atmosphere reflectances measured from geostationary satellites has been explored and leveraged for retrieving SSI estimates for more than five decades (Tarpley 1979; Müller and Pfeifroth 2022). Satellite retrievals of SSI utilize top-of-atmosphere visible and infrared reflectances to estimate the amount of solar radiation that reaches the Earth’s surface.
Data-driven solar retrieval models can outperform physical models for single ground stations (Cornejo-Bueno et al. 2019). Deep learning has enabled increases in SSI retrieval accuracy of up to 20% compared to state-of-the-art retrievals (Quesada-Ruiz et al. 2015; Hao et al. 2019; Huang et al. 2019; Verbois et al. 2023a). Recent data-driven retrieval models have been trained on pyranometer ground stations with geostationarily measured radiances as predictors to retrieve SSI by either post-processing previous SSI retrievals (Verbois et al. 2023b; Hao et al. 2020; Ma et al. 2020) or by directly estimating SSI retrievals for individual ground stations (Verbois et al. 2023a; Jiang et al. 2019; Cornejo-Bueno et al. 2019; Quesada-Ruiz et al. 2015; Hao et al. 2020; Shi et al. 2023). SSI retrievals are expected to not only achieve high accuracy and low biases within the geographic region of their training set ground stations but also well beyond. In particular, an SSI retrieval should also be accurate in regions with other surface albedos and other climates, which have not been included in the development of the SSI retrieval model (out-of-domain generalization) (Hao et al. 2020; Verbois et al. 2023a; Jiang et al. 2019; Yang et al. 2022b). Existing machine-learning based SSI retrievals were found to lack generalizability and to feature artefacts showing up in SSI estimates at locations outside their training set domains (Verbois et al. 2023a; Yang et al. 2022b).
@@
The structure of this paper is outlined as follows. Section 2 reviews existing SSI retrieval methods and concludes with the novelty of this paper. Section 3 introduces the datasets utilized in this study. The deep learning SSI retrieval and its development are detailed in section 4. Results of this study are discussed in section 5 and conclusions are provided in section 6.
@@ 2 Previous work
One of the first SSI satellite retrieval algorithms, Heliosat-1, included an empirical formula to calculate SSI (Cano et al. 1986). An empirical clear-sky model estimated SSI in the absence of clouds and assuming a given atmospheric composition. The transmissivity of the clouds was estimated by the ratio between top-of-atmosphere reflectivity measured and the minimum reflectivity over a month for cloud-free skies. The clear-sky reflectance mainly depends on the local solar zenith angle (SZA) and the surface albedo at the monitored site. Reliable surface albedo estimates are considered essential for accurate SSI retrievals (Tournadre et al. 2022). Current state-of-the-art surface solar irradiance retrieval algorithms, such as CAMS (Schroedter-Homscheidt et al. 2022), SARAH-3 (Pfeifroth et al. 2023), MSG-CPP (Roebeling et al. 2006; Greuell et al. 2013), and SolarGIS (Suri and Cebecauer 2014), use a combination of radiative transfer model (RTM) and the clear-sky index (CSI) to estimate SSI. The CSI is defined as the ratio between all-sky SSI and clear-sky SSI, , where the clear-sky SSI is calculated using a clear-sky RTM. The latter performs corrections to the clear-sky SSI due to attenuation from aerosols and trace gases. CAMS, SARAH-3 and MSG-CPP apply a cloudy-sky RTM to extract all-sky SSI based on cloud optical properties. RTM simulations are computationally expensive, so all three methods utilize look-up tables created by calculating radiative properties for various cloud and aerosol conditions. The observed cloud optical properties, aerosol properties and surface albedo are interpolated within the conditions provided by the look-up table to estimate global surface solar radiation. In MSG-CPP, the cloud optical properties are derived from near-infrared, visible light and surface albedo as parameters to a cloudy RTM look-up table (Roebeling et al. 2006; Mueller et al. 2012). While the SSI retrieval algorithms differ, they are all based on the visible and partially near-infrared reflectivity from Meteosat Second Generation satellites.
@@
Recent research on deep learning (DL) for surface solar irradiance retrieval can be categorized into two main branches: post-processing of SSI derived from radiative transfer model and direct DL retrieval models. In post-processing, DL is employed as a post-processing step for existing retrieval algorithms by using spectral imagery and derived variables as input to improve on the initial SSI retrieval estimate. Cornejo-Bueno et al. (2019) benchmarked a multitude of DL networks for post-processing Heliosat-2 and CAMS output, and found moderate improvements over a single station in Toledo, Spain. Verbois et al. (2023b) trained an extreme gradient boosting network to improve SSI estimates from HelioClim-3 (Blanc et al. 2011) based on SSI from a group of 283 pyranometers spread across France. Ma et al. (2020) inferred instantaneous values of SSI (and other radiative transfer model outputs) by training a fully-connected neural network directly on an RTM with input level L2 products from the Himawari-8 satellite. They validated the model based on ground stations across Asia and Oceania and found an average instantaneous MBE and RMSE of 8.1 and 125.9 . These methods have not been validated with regard to their generalizability to other regions outside the training set domain. Direct retrieval DL models, on the other hand, estimate SSI directly from visible or infrared satellite channels or ground station data. Quesada-Ruiz et al. (2015) trained three model ensembles of ten fully connected feed-forward neural networks each on all visible and infrared channels’ pixel values of the Spinning Enhanced Visible and Infrared Imager (SEVIRI) plus a clear-sky SSI estimate and used SSI from eight ground stations in Europe as predictands (ground truth). Each ensemble was designed to estimate hourly SSI for different weather conditions: clear-sky (CSI 0.8), intermediate (0.8 CSI 0.4), and overcast conditions (CSI 0.4). A separate ensemble network of the same number of models predicted the clear-sky index. The method achieved an average RMSE skill score improvement of 17% compared to the Heliosat-2 method (Rigollier et al. 2004) across 28 validation stations. In another study, Jiang et al. (2019) applied a convolutional residual network to estimate hourly SSI based on the Multifunctional Transport Satellites (MTSAT) series with 1616 pixel patches of visible reflectances. The network was trained on 90 pyranometer stations spread out over different land surface classes and climates in China and inferred over the entirety of China. The network performance on eight validation pyranometers within the region showed an average RMSE of 93 . They did not compare those results with an existing radiative transfer-based retrieval algorithm. Jang et al. (2022) created a CNN-based DL retrieval model trained on hourly averaged SSI measured at 81 weather stations from 2019 to 2020 across South Korea. They utilized an extra auxiliary variable the extraterrestrial solar radiation compared to (Jiang et al. 2019). The reported MBE and RMSE for the same stations from 2020 to 2021 were -1.9 and 50.0 compared to 32.8 and 90.1 for the operational RTM model. Their validation on the same training stations does not allow any conclusions made for generalization over the entire domain. Shi et al. (2023) was the first to address the importance of different level-1 satellite inputs and auxiliary inputs to direct SSI retrieval using shapely values derived from the random forest model. They trained the model to predict hourly SSI values based on visible (0.55 - 0.77), near-infrared (3.5 - 4), infrared (6.9 - 7.3, 11.5 - 12.5) from the satellite FY-4A and auxiliary variables on ground stations across China. The most important input was the solar zenith angle followed by the visible channel, solar azimuth angle, near-infrared and infrared (11.5 - 12.5) channel. Compared to (Ma et al. 2020; Jiang et al. 2019) the random forest model performed the worst with an hourly MBE and RMSE of -5.64 and 147.0 over the ground stations of China.
@@
Verbois et al. (2023a) were the first to explore the out-of-sample generalizability of DL models for SSI retrievals. They trained a fully connected neural network on the hourly SSI from pyranometers over France incorporating 13 time steps of 15 minutes patches of 33 from three SEVIRI channels: the visible and (VIS006, VIS008) and infrared (IR108). They showed the performance of the DL network to be better than CAMS with an average RMSE skill score of 20% within the training domain of southeastern France, but outside this limited domain, in the validation set over the rest of France, performance dropped off significantly to skill scores of -100% to -40% for some stations. The worse performing stations were located in coastal regions with a generally lower surface albedo which were not represented in the training dataset. Verbois et al. (2023a) analyzed if average surface albedo impacted the generalization of the DL network and found a significant Spearman correlation of 0.346 (p 0.01) between surface albedo and RMSE skill score. The results of Verbois et al. (2023a) show challenges in generalising a DL retrieval to unseen locations far differentiated from the training set.
This paper introduces an instantaneous direct retrieval model designed with two key objectives: ensuring fast inference and achieving robust generalizability across diverse and previously unseen climates and surfaces. We employ a convolutional neural network (CNN) architecture inspired by Jiang et al. (2019), utilizing only level-1 products from SEVIRI and auxiliary inputs that can be precomputed. We let it train on instantaneous SSI compared to hourly averaged SSI (Jiang et al. 2019; Quesada-Ruiz et al. 2015; Shi et al. 2023; Verbois et al. 2023b). Initially, the model is trained to emulate a radiative transfer model (RTM) using the Heliosat-3 SARAH-3 dataset. The most similar methodology was implemented by Ma et al. (2020) wherein they emulate an RTM instantaneous output but use level-2 products as inputs instead of directly retrieving the SSI from level-1.5 products here. None of the previous works have first tried to emulate an RTM and subsequently fine-tuned the model using a ground station dataset that spans only a subset of the domain. This approach aims to address the generalizability challenges of deep learning-based retrieval methods by leveraging the broad applicability of the RTM while maintaining the high performance of a direct retrieval network trained on ground stations.
@@ 3.1 Meteosat
Visible and infrared radiances from Earth that are measured by geostationary satellites provide the basis of SSI retrieval products. The Meteosat Second Generation (MSG) satellite at E, N is monitoring the regions of interest in our study – Europe and North Africa – with the Spinning Enhanced Visible and Infrared Imager (SEVIRI) that scans from East to West by rotating its sensor’s view. After each rotation, the sensor resets and scans the next line from North to South. One line scan is close to an instantaneous capture of longitudinal pixels, but individual scan lines have a delay associated with them, so an entire scan of the satellite’s disk takes approximately 12 minutes (Schmetz et al. 2002). The ground resolution of the sensor is 33 km at nadir. All eleven visible and infrared channels are included as input to the emulator model training.
3.2 Heliosat SARAH-3
Heliosat SARAH-3 (Pfeifroth et al. 2023) provides Meteosat-derived estimates of surface solar irradiance and other atmospheric variables related to clouds and radiation. It spans Meteosat First Generation (1977-2017) and MSG (2002-present) satellite retrievals from their respective instruments MVIRI and SEVIRI. The MVIRI is the predecessor of SEVIRI. SARAH-3 employs only the visible channels to keep the SARAH-3 record homogeneous in time.
The SARAH-3 algorithm draws on the cloud index method of the original Heliosat-1 algorithm by Cano et al. (1986) to derive an estimated effective cloud albedo CAL. The cloud index method calculates the ratio between measured broadband reflectance , clear-sky reflectance and maximum reflectance as CAL . The clear-sky and maximum reflectance are estimated from the frequency distribution within a month, where the maximum reflectance is taken as the quantile in the month and the clear-sky reflectance as a stationary variate of lowest values. The lowest reflectance in the month is not taken as clear-sky reflectance due to cloud shadows which give low reflectance outliers in the tail of the distribution. The effective cloud albedo is converted to a cloud index with a smooth piecewise function defined in Dagestad (2004). The observed pixel-wise cloud index serves as input to a radiative transfer look-up table to retrieve SSI (Mueller et al. 2012). Instantaneous SSI from SARAH-3 is published with a two-week delay at a frequency of 30 minutes, so every second SEVIRI capture.